OpenStudy (anonymous):

sheri bob and mark can inspect 40 disk drives per hour when all three are working. sheri and bob working together can inspect 28 per hour, and sheri and mark together can inspect 27 per hour. how many can each person inspect per hour?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

S+b+m=40 s+b=28 s+m=27 Equation 1-equation 2 M= 40-28=12 s=15 b=13

OpenStudy (anonymous):

S means sheri M means Mark B means bob

OpenStudy (anonymous):

thank you so much for your help!!